candice rakes leaves for her neighbors. She charges $4 per hour to rake leaves plus $7 to remove them. She uses the equation t=7+4h. How much will candice earn if she works for 3 hours and 15 minutes

Answer by fcabanski(1391)

15 minutes is .25 hour, so 3 hours and 15 minutes is 3.25 hours = h.


t = 7 + 4*3.5 = 7 + 14 = $21

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)